# Beyond the Use-and-then-Forget (UatF) Bound: Fixed Point Algorithms for Statistical Max-Min Power Control

### v2

@@ I Introduction
Against this background, we demonstrate that, in networks where the beamforming strategy is independent of the transmit power vector (the general case will be considered elsewhere), alternatives to the UatF bound – e.g., the coherent decoding bound and the upper bound in [8, Proposition 1] – can be employed in tractable weighted max-min optimization problems, provided we accept that expectations are approximated via Monte Carlo sampling, as also done in power control problems based on the UatF bound [2, 8, 7, 9]. Nevertheless, we avoid additional approximations, such as replacing the expectation of a ratio with the ratio of expectations.
To further contextualize our contribution, we note that many existing power control algorithms largely fall into two classes: (i) per-sample schemes, which solve an optimization problem for each channel sample [10, 11], but suffer from poor scalability in distributed MIMO systems; and (ii) statistical-level schemes, which solve a single optimization problem per channel distribution [1, 2], but rely on the UatF bound and thus inherit its limitations. Our framework builds upon the foundations of both classes to address these drawbacks. In particular, we use this framework to derive a fixed point algorithm – built on the iterations introduced to the wireless literature in [11] – that provably converges to globally optimal solutions of (weighted) max-min problems using an achievable rate bound that exploits instantaneous CSI at the decoder.
@@ II-A Notation and definitions
The following technical result can be established with arguments analogous to those in [12, Lemmas 2 and 3][13], in combination with the standard result shown in [14, Theorem 10.3][15, Proposition 1].
@@
A probability space is denoted by $(\Omega,\mathcal{F},\mathsf{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ is the event space, and $\mathsf{P}$ is the probability measure. The next result is a direct implication of [16, Proposition 2.16].
@@
Given a probability space $(\Omega,\mathcal{F},\mathsf{P})$ and a scalar $\delta\geq 0$, assume that a random variable $X:\Omega\to{\mathbb{R}}$ satisfies $X>\delta$ almost surely (a.s.) (or, equivalently, $X(\omega)>\delta$ for almost every $\omega\in\Omega$) and $E[X(\omega)]:=\int X(\omega)\mathsf{P}(\mathrm{d}\omega)$ is well-defined (i.e., $E[X(\omega)]\in{\mathbb{R}}_{+}$). [^1] We adopt the stochastic-programming notation $E[X(\omega)]$ rather than $E[X]$ to avoid ambiguity when deterministic decision variables are present. This distinction is especially important later in (5) and (8), where the notation makes explicit that the power vector is not a decision variable for the beamformers, which are random vectors. Then $E[X(\omega)]>\delta$.
@@ II-B Framework for Power Control with Tight Ergodic Rate Bounds
The key distinction between Definition 1 and the original standard interference functions in [10] lies in the domain of the mappings. Standard interference functions restricted to the positive cone ${\mathbb{R}}_{++}^{N}$ are MSP functions by definition, and both classes of functions are continuous on ${\mathbb{R}}_{++}^{N}$ [17]. However, the continuous extensions of MSP functions to the nonnegative cone ${\mathbb{R}}_{+}^{N}$, which are guaranteed to exist [17][18, Theorem 5.1.5], are not standard interference functions in general. Without additional assumptions, these existing results and the positivity property in Definition 1 guarantee only the preservation of positivity, monotonicity, and weak scalability on ${\mathbb{R}}_{+}^{N}\setminus{\mathbb{R}}_{++}^{N}$– i.e., the strict inequality in the scalability condition has to be relaxed to a weak inequality in general. For example, consider the continuous extension $\bar{f}:{\mathbb{R}}^{2}_{+}\to{\mathbb{R}}_{++}$ of the function $f:{\mathbb{R}}^{2}_{++}\to{\mathbb{R}}_{++}:(x_{1},x_{2})\mapsto 1+x_{1}+\sqrt{x_{2}}+\max\{x_{1}-1,0\}$ to the domain ${\mathbb{R}}_{+}^{2}$, which is just the function $f$ itself with its domain extended to ${\mathbb{R}}_{+}^{2}$. It satisfies all properties in Definition 1 on ${\mathbb{R}}_{++}^{2}$, but, for ${\bm{x}}=(1,0)\notin{\mathbb{R}}_{++}^{2}$, we do not have scalability because $(\forall\alpha>1)~\bar{f}(\alpha{\bm{x}})=\alpha\bar{f}({\bm{x}})$.
@@
Although the above aspects might prevent the direct application of some results for standard interference functions to MSP functions – because, as discussed above, the set of continuous standard interference functions restricted to the positive cone is a proper subset of MSP functions [^2] In turn, MSP functions are a proper subclass of order-preserving subhomogeneous functions [18, 20]. –, many key properties remain analogous. In particular, the class of MSP functions is closed under addition, positive scaling, and finite maxima and minima, just as in the case of standard interference functions. Furthermore, any MSP mapping $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ is contractive in the complete metric space $({\mathbb{R}}_{++}^{N},d)$ [18, Lemma 2.1.7] (but not necessarily a Lipschitz contraction), where $d:{\mathbb{R}}_{++}^{N}\times{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}$ is Thompson’s metric defined in [18, p. 30], so $T$ has at most one fixed point. If $\mathrm{Fix}(T)\neq\emptyset$, arguments similar to those described in [10] to prove convergence of the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via the fixed point iterations $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=T({\bm{x}}_{n})$, with ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, can be used for MSP mappings. In addition, the following property, which is crucial for the power control algorithms derived in the next section, can be established using the same arguments in [11], and they are therefore omitted.
@@
Furthermore, the solution $(\gamma^{\star},{\bm{x}}^{\star})$ to (1) (called the conditional eigenpair) is such that the conditional eigenvalue $\gamma^{\star}$ satisfies $\gamma^{\star}=\|T({\bm{x}}^{\star})\|$ and the conditional eigenvector $\bm{x}$$^{\star}$ is the unique fixed point ${\bm{x}}^{\star}\in\mathrm{Fix}(G)$ of the mapping $G:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(1/\|T({\bm{x}})\|)T({\bm{x}})$. In addition, given ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=G({\bm{x}}_{n})$ converges to $\bm{x}$$^{\star}$.
@@
$z:{\mathbb{R}}_{++}^{N}\times\Omega\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto h({\bm{x}})/g({\bm{x}},\omega),$
is an MSP function. Further assume that there exists a random variable $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$, we have $z({\bm{x}},\omega)\geq\delta(\omega)>0$ for almost every $\omega\in\Omega$, and that the expectations $E[(\delta(\omega))^{-1}]$ and $E[g({\bm{x}},\omega)]:=\int g({\bm{x}},\omega)~\mathsf{P}(\mathrm{d}\omega)\geq 0$ are well defined for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ (i.e., $E[(\delta(\omega))^{-1}]\in{\mathbb{R}}_{++}$ and $(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~E[g({\bm{x}},\omega)]\in{\mathbb{R}}_{+}$). Then
@@
(Positivity) For almost every $\omega\in\Omega$, the function $q_{\omega}$ is an MSP function by assumption, so it can be continuously extended to the domain ${\mathbb{R}}_{+}^{N}$ while preserving monotonicity [18, Theorem 5.1.5]. We denote this continuous extension of $q_{\omega}$ to the domain ${\mathbb{R}}_{+}^{N}$ by $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$. Let $({\bm{y}}_{n})_{n\in{\mathbb{N}}}$ be a sequence in ${\mathbb{R}}_{++}^{N}$ converging to $\bm{0}$}. Hence, from the assumptions of the proposition, there exists $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for almost every $\omega\in\Omega$, we have
@@ III Power control in multi-user MIMO
We consider the uplink of a general multi-user MIMO network with $L\in{\mathbb{N}}$ access points, each equipped with $M\in{\mathbb{N}}$ antennas. Thus, the total number of antennas on the access point side is $K:=LM$. The set of access points is denoted by $\mathcal{A}:=\{1,\ldots,L\}$. The system also includes $N\in{\mathbb{N}}$ single-antenna users, which are indexed by the set $\mathcal{U}:=\{1,\ldots,N\}$. The channel and beamforming vectors for user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ are random (column) vectors denoted by, respectively, ${\bm{h}}_{u,a}:\Omega\to\mathbb{C}^{M}$ and ${\bm{v}}_{u,a}:\Omega\to\mathbb{C}^{M}$, where $(\Omega,\mathcal{F},\mathsf{P})$ is the probability space. The transmit power of user $u\in\mathcal{U}$ is denoted by $p_{u}\in{\mathbb{R}}_{++}$. Without any loss of generality, we assume that $(\forall\omega\in\Omega)~{\bm{v}}_{u,a}(\omega)={\bm{0}}$ if access point $a\in\mathcal{A}$ does not participate in the processing of the signal transmitted by user $u\in\mathcal{U}$, and that $\|{\bm{v}}_{u}(\omega)\|_{2}=1$ for almost every $\omega\in\Omega$ and every $u\in\mathcal{U}$. The aggregated channel vector and aggregated beamforming vector for user $u\in\mathcal{U}$ are denoted by, respectively, ${\bm{h}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{h}}_{u,1}^{t}(\omega),\ldots,{\bm{h}}^{t}_{u,L}(\omega)]^{t}$ and ${\bm{v}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{v}}^{t}_{u,1}(\omega),\ldots,{\bm{v}}^{t}_{u,L}(\omega)]^{t}$.
The above model covers many common technologies in the literature, including multi-cell massive MIMO networks as defined in [3] and user-centric cell-less networks with different levels of cooperation as defined in [2]. For brevity and notational simplicity (see Remark 3 below), we evaluate the utility (performance) of user $u\in\mathcal{U}$ as a function of the power allocation ${\bm{p}}=[p_{1},\ldots,p_{N}]^{t}\in{\mathbb{R}}_{++}^{N}$ via the following well-known achievable rate under perfect channel state information (CSI) [8, Proposition 1][21, Lemma 1]:
@@
The utility in (5) is often called the “genie-aided” or “optimistic” upper bound for the achievable ergodic rates (when interference is treated as noise) because it assumes perfect CSI at the receiver [2, p. 328]. Nevertheless, the results we derive can be readily extended to other bounds, such as the coherent decoding lower bound in [2, Theorem 5.1][3, Theorem 4.1]. While this bound uses channel estimates – instead of perfect channels – and introduces an “error covariance” term in the denominator of the expectation in (5), the subsequent analysis in this study applies to it in a line-by-line manner with only straightforward changes.
@@
For every user $u\in\mathcal{U}$ and power vector ${\bm{p}}\in{\mathbb{R}}^{N}_{++}$, assume that the expectation in (5) is well defined, $E[|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}]<\infty$, and $|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|\neq 0$ for almost every $\omega\in\Omega$. Then $f_{u}:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}$ in (9) is an MSP function.
@@
is concave $s_{u}$ is defined in (8)], and it has a continuous concave extension $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}$ satisfying $(\forall{\bm{p}}\in{\mathbb{R}}_{++}^{N})~q_{\omega}({\bm{p}})\geq\bar{q}_{\omega}({\bm{0}})=\sigma^{2}/|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}>0$. Now apply [15, Proposition 1] to $\bar{q}_{\omega}$ to conclude that ${q}_{\omega}$, the restriction of $\bar{q}_{\omega}$ to the domain ${\mathbb{R}}_{++}^{N}$, is an MSP function. The desired result now follows as an immediate application of Proposition 2. ∎
We now have all the necessary background to pose and solve a power control problem that has the objective of maximizing the minimum (weighted) achievable rate while satisfying the power constraint $\|{\bm{p}}\|\leq p_{\mathrm{max}}$, where $\|\cdot\|$ is a given monotone norm (typically the $l_{\infty}$-norm in the uplink) and $p_{\mathrm{max}}>0$ the maximum transmit power. Formally, we consider the following optimization problem:
@@
We can use the same arguments in [5, Proposition 2] with standard interference functions replaced with MSP functions to show that, although the solution to (16) (and, hence, (12)) may not be unique, there always exists a solution achieving equality in all constraints; i.e., there exists a solution $(\gamma^{\star},{\bm{p}}^{\star})$ satisfying $(\gamma^{\star})^{-1}{\bm{p}}^{\star}=T({\bm{p}}^{\star})$ and $(1/p_{\text{max}})\|{\bm{p}}^{\star}\|=1$. These equalities establish that $(\gamma^{\star})^{-1}$ and $\bm{p}$$^{\star}$ are, respectively, the conditional eigenvalue and conditional eigenvector of the mapping $T$ and the monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|_{\star}:=(1/p_{\text{max}})\|{\bm{p}}\|$. Therefore, as a corollary of Proposition 1, Proposition 3, and [5, Proposition 2] (after adapting the proof of this last proposition by substituting standard interference functions with MSP functions), we obtain the following result, which shows a simple fixed point method that provably converges to the solution to (12) with minimum total power:
@@
$\text{minimize}_{{\bm{p}}\in\mathcal{S}}\|{\bm{p}}\|_{1}.$

### v3

@@ I Introduction
Against this background, we demonstrate that, in networks where the beamforming strategy is independent of the transmit power vector (the general case will be considered elsewhere), alternatives to the UatF bound – e.g., the coherent decoding bound and the upper bound in [18, Proposition 1] – can be employed in tractable weighted max-min optimization problems, provided we accept that expectations are approximated via Monte Carlo sampling, as also done in power control problems based on the UatF bound [11, 18, 2, 20]. Nevertheless, we avoid additional approximations, such as replacing the expectation of a ratio with the ratio of expectations.
To further contextualize our contribution, we note that many existing power control algorithms largely fall into two classes: (i) per-sample schemes, which solve an optimization problem for each channel sample [23, 21], but suffer from poor scalability in distributed MIMO systems; and (ii) statistical-level schemes, which solve a single optimization problem per channel distribution [19, 11], but rely on the UatF bound and thus inherit its limitations. Our framework builds upon the foundations of both classes to address these drawbacks. In particular, we use this framework to derive a fixed point algorithm – built on the iterations introduced to the wireless literature in [21] – that provably converges to globally optimal solutions of (weighted) max-min problems using an achievable rate bound that exploits instantaneous channel state information (CSI) at the decoder.
@@ II-A Notation and definitions
The following technical result can be established with arguments analogous to those in [7, Lemmas 2 and 3][5], in combination with the standard result shown in [22, Theorem 10.3][9, Proposition 1].
@@
A probability space is denoted by $(\Omega,\mathcal{F},\mathsf{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ is the event space, and $\mathsf{P}$ is the probability measure. The next result is a direct implication of [12, Proposition 2.16].
@@
Given a probability space $(\Omega,\mathcal{F},\mathsf{P})$ and a scalar $\delta\geq 0$, assume that a random variable $X:\Omega\to[0,\infty[$ satisfies $X>\delta$ ($\mathsf{P}$-)almost surely (a.s.) (or, equivalently, $X(\omega)>\delta$ for almost every $\omega\in\Omega$) and $E[X(\omega)]:=\int X(\omega)\mathsf{P}(\mathrm{d}\omega)$ is well-defined (i.e., $E[X(\omega)]\in{\mathbb{R}}_{+}$). [^1] We adopt the stochastic-programming notation $E[X(\omega)]$ rather than $E[X]$ to avoid ambiguity when deterministic decision variables are present. This distinction is especially important later in (5) and (8), where the notation makes explicit that the power vector is not a decision variable for the beamformers, which are random vectors. Then $E[X(\omega)]>\delta$.
@@ II-B Framework for Power Control with Tight Ergodic Rate Bounds
The key distinction between Definition 1 and the original standard interference functions in [23] lies in the domain of the mappings. Standard interference functions restricted to the positive cone ${\mathbb{R}}_{++}^{N}$ are MSP functions by definition, and both classes of functions are continuous on ${\mathbb{R}}_{++}^{N}$ [4]. However, the continuous extensions of MSP functions to the nonnegative cone ${\mathbb{R}}_{+}^{N}$, which are guaranteed to exist [4][14, Theorem 5.1.5], are not standard interference functions in general. Without additional assumptions, these existing results and the positivity property in Definition 1 guarantee only the preservation of positivity, monotonicity, and weak scalability on ${\mathbb{R}}_{+}^{N}\setminus{\mathbb{R}}_{++}^{N}$– i.e., the strict inequality in the scalability condition has to be relaxed to a weak inequality in general. For example, consider the continuous extension $\bar{f}:{\mathbb{R}}^{2}_{+}\to{\mathbb{R}}_{++}$ of the function $f:{\mathbb{R}}^{2}_{++}\to{\mathbb{R}}_{++}:(x_{1},x_{2})\mapsto 1+x_{1}+\sqrt{x_{2}}+\max\{x_{1}-1,0\}$ to the domain ${\mathbb{R}}_{+}^{2}$, which is just the function $f$ itself with its domain extended to ${\mathbb{R}}_{+}^{2}$. It satisfies all properties in Definition 1 on ${\mathbb{R}}_{++}^{2}$, but, for ${\bm{x}}=(1,0)\notin{\mathbb{R}}_{++}^{2}$, we do not have scalability because $(\forall\alpha>1)~\bar{f}(\alpha{\bm{x}})=\alpha\bar{f}({\bm{x}})$.
@@
Although the above aspects might prevent the direct application of some results for standard interference functions to MSP functions – because, as discussed above, the set of standard interference functions restricted to the positive cone is a proper subset of MSP functions [^2] In turn, MSP functions are a proper subclass of order-preserving subhomogeneous functions [14, 13]. –, many key properties remain analogous. In particular, the class of MSP functions is closed under addition, positive scaling, and finite maxima and minima, just as in the case of standard interference functions. Furthermore, any MSP mapping $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ is contractive in the complete metric space $({\mathbb{R}}_{++}^{N},d)$ [14, Lemma 2.1.7] (but not necessarily a Lipschitz contraction), where $d:{\mathbb{R}}_{++}^{N}\times{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}$ is Thompson’s metric defined in [14, p. 30], so $T$ has at most one fixed point. If $\mathrm{Fix}(T)\neq\emptyset$, arguments similar to those described in [23] to prove convergence of the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via the fixed point iterations $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=T({\bm{x}}_{n})$, with ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, can be used for MSP mappings. In addition, the following property, which is crucial for the power control algorithms derived in the next section, can be established using the same arguments in [21], and they are therefore omitted.
@@
Furthermore, the solution $(\gamma^{\star},{\bm{x}}^{\star})$ to (1) (called the conditional eigenpair) is such that the conditional eigenvalue $\gamma^{\star}$ satisfies $\gamma^{\star}=\|T({\bm{x}}^{\star})\|$, and the conditional eigenvector $\bm{x}$$^{\star}$ is the unique fixed point ${\bm{x}}^{\star}\in\mathrm{Fix}(G)$ of the mapping $G:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(1/\|T({\bm{x}})\|)T({\bm{x}})$. In addition, given ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=G({\bm{x}}_{n})$ converges to $\bm{x}$$^{\star}$.
@@
$z:{\mathbb{R}}_{++}^{N}\times\Omega\to{\mathbb{R}}_{++}:({\bm{x}},\omega)\mapsto h({\bm{x}})/g({\bm{x}},\omega),$
is an MSP function. Further assume that there exists a random variable $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ and for almost every $\omega\in\Omega$, we have $z({\bm{x}},\omega)\geq\delta(\omega)>0$, and that the expectations $E[(\delta(\omega))^{-1}]$ and $E[g({\bm{x}},\omega)]:=\int g({\bm{x}},\omega)~\mathsf{P}(\mathrm{d}\omega)\geq 0$ are well defined for every ${\bm{x}}\in{\mathbb{R}}_{++}^{N}$ (hence $E[(\delta(\omega))^{-1}]\in{\mathbb{R}}_{++}$ and $(\forall{\bm{x}}\in{\mathbb{R}}_{++}^{N})~E[g({\bm{x}},\omega)]\in{\mathbb{R}}_{+}$). Then
@@
(Positivity) For almost every $\omega\in\Omega$, the function $q_{\omega}$ is an MSP function by assumption, so it can be continuously extended to the domain ${\mathbb{R}}_{+}^{N}$ while preserving monotonicity [14, Theorem 5.1.5]. We denote this continuous extension of $q_{\omega}$ to the domain ${\mathbb{R}}_{+}^{N}$ by $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$. Let $({\bm{y}}_{n})_{n\in{\mathbb{N}}}$ be a sequence in ${\mathbb{R}}_{++}^{N}$ converging to $\bm{0}$}. Hence, from the assumptions of the proposition, there exists $\delta:\Omega\to{\mathbb{R}}_{++}$ such that, for almost every $\omega\in\Omega$, we have
@@ III Power control in multi-user MIMO
We consider the uplink of a general multi-user MIMO network with $L\in{\mathbb{N}}$ access points, each equipped with $M\in{\mathbb{N}}$ antennas. Thus, the total number of antennas on the access point side is $K:=LM$. The set of access points is denoted by $\mathcal{A}:=\{1,\ldots,L\}$. The system also includes $N\in{\mathbb{N}}$ single-antenna users, which are indexed by the set $\mathcal{U}:=\{1,\ldots,N\}$. The channel and beamforming vectors for user $u\in\mathcal{U}$ and access point $a\in\mathcal{A}$ are random (column) vectors denoted by, respectively, ${\bm{h}}_{u,a}:\Omega\to\mathbb{C}^{M}$ and ${\bm{v}}_{u,a}:\Omega\to\mathbb{C}^{M}$, where $(\Omega,\mathcal{F},\mathsf{P})$ is the probability space as defined below Fact 1. The transmit power of user $u\in\mathcal{U}$ is denoted by $p_{u}\in{\mathbb{R}}_{++}$. Without any loss of generality, we assume that $(\forall\omega\in\Omega)~{\bm{v}}_{u,a}(\omega)={\bm{0}}$ if access point $a\in\mathcal{A}$ does not participate in the processing of the signal transmitted by user $u\in\mathcal{U}$, and that $\|{\bm{v}}_{u}(\omega)\|_{2}=1$ for almost every $\omega\in\Omega$ and every $u\in\mathcal{U}$. The aggregated channel vector and aggregated beamforming vector for user $u\in\mathcal{U}$ are denoted by, respectively, ${\bm{h}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{h}}_{u,1}^{t}(\omega),\ldots,{\bm{h}}^{t}_{u,L}(\omega)]^{t}$ and ${\bm{v}}_{u}:\Omega\to\mathbb{C}^{K}:\omega\mapsto[{\bm{v}}^{t}_{u,1}(\omega),\ldots,{\bm{v}}^{t}_{u,L}(\omega)]^{t}$.
The above model covers many common technologies in the literature, including multi-cell massive MIMO networks as defined in [3] and user-centric cell-less networks with different levels of cooperation as defined in [11]. For brevity and notational simplicity (see Remark 3 below), we evaluate the utility (performance) of user $u\in\mathcal{U}$ as a function of the power allocation ${\bm{p}}=[p_{1},\ldots,p_{N}]^{t}\in{\mathbb{R}}_{++}^{N}$ via the following well-known achievable rate under perfect CSI [18, Proposition 1][6, Lemma 1]:
@@
The utility in (5) is often called the “genie-aided” or “optimistic” upper bound for the achievable ergodic rates (when interference is treated as noise) because it assumes perfect CSI at the receiver [11, p. 328]. Nevertheless, the results we derive can be readily extended to other bounds, such as the coherent decoding lower bound in [11, Theorem 5.1][3, Theorem 4.1]. While this bound uses channel estimates – instead of perfect channels – and introduces an “error covariance” term in the denominator of the expectation in (5), the subsequent analysis in this study applies to it in a line-by-line manner with only straightforward changes.
@@
For every user $u\in\mathcal{U}$ and power vector ${\bm{p}}\in{\mathbb{R}}^{N}_{++}$, assume that the expectation in (5) is well defined, $E[|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}]<\infty$, and $|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|\neq 0$ for almost every $\omega\in\Omega$. Then $f_{u}:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{++}$ in (9) is an MSP function for every $u\in\mathcal{U}$.
@@
is concave $s_{u}$ is defined in (8)], and it has a continuous concave extension $\bar{q}_{\omega}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}$ satisfying $(\forall{\bm{p}}\in{\mathbb{R}}_{++}^{N})~q_{\omega}({\bm{p}})\geq\bar{q}_{\omega}({\bm{0}})=\sigma^{2}/|{\bm{h}}_{u}(\omega)^{H}{\bm{v}}_{u}(\omega)|^{2}>0$. Now apply [9, Proposition 1] to $\bar{q}_{\omega}$ to conclude that ${q}_{\omega}$, the restriction of $\bar{q}_{\omega}$ to the domain ${\mathbb{R}}_{++}^{N}$, is an MSP function. The desired result now follows as an immediate application of Proposition 2. ∎
We now have all the necessary background to pose and solve a power control problem that has the objective of maximizing the minimum (weighted) achievable rate while satisfying the power constraint $\{{\bm{p}}\in{\mathbb{R}}_{++}^{N}\mid\|{\bm{p}}\|\leq p_{\mathrm{max}}\}$, where $\|\cdot\|$ is a given monotone norm (typically the $l_{\infty}$-norm in the uplink) and $p_{\mathrm{max}}>0$ the maximum transmit power. Formally, we consider the following optimization problem:
@@
We can use the same arguments in [8, Proposition 2] with standard interference functions replaced with MSP functions to show that, although the solution to (16) (and, hence, (12)) may not be unique, there always exists a solution achieving equality in all constraints; i.e., there exists a solution $(\gamma^{\star},{\bm{p}}^{\star})$ satisfying $(\gamma^{\star})^{-1}{\bm{p}}^{\star}=T({\bm{p}}^{\star})$ and $(1/p_{\text{max}})\|{\bm{p}}^{\star}\|=1$. These equalities establish that $(\gamma^{\star})^{-1}$ and $\bm{p}$$^{\star}$ are, respectively, the conditional eigenvalue and conditional eigenvector of the mapping $T$ and the monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|_{\star}:=(1/p_{\text{max}})\|{\bm{p}}\|$. Therefore, as a corollary of Proposition 1, Proposition 3, and [8, Proposition 2] (after adapting the proof of this last proposition by substituting standard interference functions with MSP functions), we obtain the following result, which shows a simple fixed point method that provably converges to the solution to (12) with minimum total power:
@@
$\mathrm{minimize}_{{\bm{p}}\in\mathcal{S}}\|{\bm{p}}\|_{1}.$